OpenStudy (anonymous):

please explain. Solve the equation using the quadratic formula. 6/x+6/x+9=1

jhonyy9 (jhonyy9):

6 6 --- + ----- = 1 so the common denominator will be x(x+9) x x+9 6(x+9) +6x ------------ = 1(x)(x+9) / (x(x+9) x(x+9) 6x+54 +6x = x(x+9) 12x +54 = x^2 +9x x^2 +9x -12x -54 =0 x^2 -3x -54 =0 (x-9)(x+6) =0 x-9=0 x=9 x+6=0 x=-6 hope this will help you